Can someone show me how to solve the following linear story problem?

Christopher orders a 3 topping pizza for $15.35,and a 5 topping pizza for $17.95. Write and solve a system of linear equations to find the price of a plain cheese pizza (no toppings) and the cost of each topping. I sure need help on this.
Thanks a lot.

Found 3 solutions by jorel555, MathTherapy, solver91311:
Answer by jorel555(1290)

(Edited:12-1-16)
Let m be the cost of a plain cheese pizza, and n be the price for each extra topping. Then:
m+3n=15.35
m+5n=17.95
Subtracting the first from the second, we get:
2n=2.60
n=1.3
Toppings are $1.30 each; and a plain cheese pizza costs $11.45 ☺☺☺☺

Answer by MathTherapy(10552)


Can someone show me how to solve the following linear story problem?

Christopher orders a 3 topping pizza for $15.35,and a 5 topping pizza for $17.95. Write and solve a system of linear equations to find the price of a plain cheese pizza (no toppings) and the cost of each topping. I sure need help on this.
Thanks a lot.
Let cost of a plain pizza be P, and cost of each topping, T
Then we get: P + 3T = 15.35 ------ eq (i)
Also, P + 5T = 17.95 ------- eq (ii)
- 2T = - 2.6 ------- Subtract eq (ii) from eq (i)
T, or
P + 3(1.3) = 15.35 ------- Substituting 1.3 for T in eq (i)
P + 3.9 = 15.35
P, or

Answer by solver91311(24713)



Let represent the cost of a plain cheese pizza. Let represent the cost of one extra topping. Then:





Solve the system for and

John

My calculator said it, I believe it, that settles it